Annual demand for educational toys in the city of Lambert is 1,200 units. The fixed ordering...

  1. Annual demand for educational toys in the city of Lambert is 1,200 units. The fixed ordering cost is $41, and the holding cost is known to be always $2 per unit per year regardless of unit price. The contract offered by the supplier is an all-units discount and consists of the following price schedule:

Quantity

Unit Price

1 to 199

$ 27

200 to 299

$ 26

300 to 399

$ 25

400 or more

$ 24

  1. What are the realizable EOQ(s)?

Answer:

  1. If the largest realizable EOQ quantity in i) is ordered, what is the corresponding unit price applied by the supplier?

Answer:

  1. Compare the total cost for the largest realizable EOQ in part i) with the total costs for all larger break points. What is the optimal order quantity?

Answer:
Answer #1

DEMAND = 1200

ORDERING COST = 41

HOLDING COST = 2

EOQ = SQRT(2 * DEMAND * ORDERING COST / HOLDING COST) = SQRT(2 * 1200 * 41 / 2) = 222

ANNUAL HOLDING COST = ADJUSTED EOQ / 2 * HOLDING COST PER UNIT

ANNUAL ORDERING COST = (ANNUAL DEMAND / ADJUSTED EOQ) * ORDERING COST

ANNUAL MATERIAL COST = ANNUAL DEMAND * OFFERED PRICE PER UNIT

TOTAL COST OF INVENTORY = ANNUAL(HOLDING + ORDERING + MATERIAL)

OPTIMAL ORDER QUANTITY = 400

ASSOCIATED COST = 29323


NO.

LOWER LIMIT

UPPER LIMIT

PER UNIT PRICE

HOLDING COST

EOQ

ADJUSTED QUANTITY

ANNUAL HOLDING COST

ANNUAL ORDER COST

ANNUAL MATERIAL COST

TOTCAL COST OF INVENTORY

1

0

199

27

2

222

199

199 / 2 = 199

1200 / 199 * 41 = 247.24

1200 * 27 = 32400

199 + 247.24 + 32400 = 32846

2

200

299

26

2

222

222

222 / 2 = 222

1200 / 222 * 41 = 221.62

1200 * 26 = 31200

222 + 221.62 + 31200 = 31644

3

300

399

25

2

222

300

300 / 2 = 300

1200 / 300 * 41 = 164

1200 * 25 = 30000

300 + 164 + 30000 = 30464

4

400

OR MORE

24

2

222

400

400 / 2 = 400

1200 / 400 * 41 = 123

1200 * 24 = 28800

400 + 123 + 28800 = 29323